OpenStudy (anonymous):

A card is selected from a standard deck of 52 playing cards. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards). Find the probability of selecting · a five given the card is a not a club. · a heart given the card is red. · a face card, given that the card is black. Show step by step work. Give all solutions exactly in reduced fraction form. HELP PLEASE

OpenStudy (anonymous):

1) 1/4 of the deck are clubs which is 3 face cards. 12-3=9. So, you have 9/52 chance of selecting a face card. 2) I believe there are 2 colors, red and black, right? If so, then 1/2 are red. 52/2 = 26 3) 1/2 of the cards are black, so 12/2= 6. You have a 6/52 chance of pulling a black face card.

OpenStudy (anonymous):

THX